OpenStudy (anonymous):

Mary wants to buy one large pizza, one medium pizza, and three drinks. The drinks cost $p$ dollars each, the medium pizza costs two times as much as one drink, and the large pizza costs three times as much as one drink. If Mary started with $30$ dollars, how much money would she have left after making all of her purchases? Give your answer as an expression of $p$.

OpenStudy (xapproachesinfinity):

well start like this, 3 drinks for p dollars each so 3p dollars for all of the three drinks then the medium pizza costs 2 times as much as one drink so 2p dollars for one medium pizza then the large pizza costs 3p dollars so 30 -(3p +2p +3p)=30-8p

OpenStudy (xapproachesinfinity):

that's the expression you are looking for

OpenStudy (anonymous):

ok thanks!

OpenStudy (xapproachesinfinity):

if they give us what p is we would know how much each one costs

OpenStudy (xapproachesinfinity):

let's assume that one drink costs 1,25 dollars how much is the cost of large pizza and medium pizza?

OpenStudy (anonymous):

I don't get what the answer is

OpenStudy (xapproachesinfinity):

well look again and read the info one large pizza cost 3 times as much as the cost of one drink we assumed one drink costs 1,25 so 1,25 +1,25+1,25=$3,75 medium costs twice as much as the cost of one drink so 1,25+1,25=2,50 three drinks would be 3,75 so the total that Mary will be charged is 3,75+3,75+2,50=$10

OpenStudy (xapproachesinfinity):

i'm just giving you the sense of what's going on using an example

OpenStudy (xapproachesinfinity):

this is something that you do in your everyday life

OpenStudy (anonymous):

oh I got the answer right thanks

OpenStudy (xapproachesinfinity):

Weclcome